OpenStudy (anonymous):

What are the approximate solutions of 4x^2 + 3 = −12x to the nearest hundredth? x ≈ −3.23 and x ≈ 0.23 x ≈ −2.72 and x ≈ −0.28 x ≈ 0.28 and x ≈ 2.72 x ≈ −0.23 and x ≈ 3.23

OpenStudy (anonymous):

I dont get this AT ALL :/

OpenStudy (anonymous):

Rearrange to form a quadratic: 4x^2 + 3 = -12 4x^2 + 3 + 12x = 0 4x^2 + 12x + 3 = 0 Then solve using the quadratic equation, where the quadratic is of the form ax^2 + bx + c = 0: \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (campbell_st):

the equation is \[4x^2 + 12x + 3 = 0\] you need the general quadratic formula with a = 4, b = 12 and c = 3 \[x = \frac{-12\pm \sqrt{12^2 - 4 \times4 \times3}}{2\times4}\] the solutions are \[x = \frac{ -12 p \sqrt{96}}{8}...and...x = \frac{-12 - \sqrt{96}}{8}\] just evaluate for your answers

OpenStudy (campbell_st):

oops \[x = \frac{-12 + \sqrt{96}}{8}\]

OpenStudy (anonymous):

so just evaluate from there and thats the answer?

OpenStudy (campbell_st):

thats correct.... givwn both values of x are negative... it should be easy to find the solutions in your choices.